In a two-good model, a consumer starts with an intial endowment of 12.00 units of good 1 and 10.00 units of good 2. If this consumer sold all of her units of good 2, she would be able to afford 35.00 additional units of good 1. Given this information, what is the slope of this consumer's budget line?


1
Expert's answer
2021-10-18T11:27:15-0400

Solution:

The slope of a consumer’s budget line ="-(\frac{Price\; of\; Good \; 1}{Price\; of\; Good \; 2} ) or (\frac{Quantity\; of\; Good \; 1}{Price\; of\; Good\; 2} )"


Since we only have the quantities, we will use the quantity formula:


For every 10 units sold of Good 2, she would afford 35 units of Good 1


Budget line slope = "\frac{35}{10} = 3.5"